Maria loaned both of her children $10,000 interest free on January 1, 2019. Bobby used his loan proceeds to make a down payment on his personal residence. His net investment income was $1,025 for 2019. Shelia used her loan proceeds to purchase Amazon stock. Her net investment income was $1,010 for 2019. Assuming Maria was not attempting to shift income to the children, and the applicable Federal rate is 10%, how much interest income must she include on her 2019 return? The loans were outstanding for the entire year with interest compounded semiannually.

$1,010

$1,025

$2,050

$1,000
Answer #1

Answer is option B

B. $1,025

Interest =amount – prinicipal

Amount = principal *(1+r)^n = (10000*(1.05^2)) = 11025

R = 10%/2 = 5%

N= 1*2 = 2

Interest = 11025 -10000 = $1025